# Generalized Einstein-Rosen bridge inside black holes

### v2

@@ 4 Symmetry and asymmetry of test particle trajectories near rotating black hole
Figs. 57 demonstrate both symmetry and asymmetry features of massive and massless particle trajectories plunging into rotating Kerr black hole with spin $a=0.9982$. Magenta arrows show the direction of the black hole rotation in accordance with the gimlet rule. Multicolored curves at Figs. 7 and 8 are the geodesic trajectories for massless test particles like photons ($\mu=0$) numerically calculated with using equations of motion (3)–(6). By approaching the black hole, the trajectories of all particles, both massive and massless ones, are infinitely winding up on the black hole event horizon in the direction of black hole rotation and at the fixed latitudes. This winding up is a symmetry manifestation of all trajectories behavior, which are plunging into rotating black hole. At the same direction of black hole rotation is a corresponding asymmetry manifestation of the gravitational field of the Kerr metric.
At last, for completeness of black hole symmetric and asymmetric properties, at Figs. 8 is shown the trajectory of the test planet ($\mu\neq 0$) with parameters $\gamma=1$, $\lambda=1$ and $Q=0.5$. This test planet is plunging into the spherically symmetric and nonrotating Schwarzschild black hole (with both the spin $a=0$ and electric charge $q=0$), starting from the radial distance $r=6$.
It must be especially checked that the traversable (though only one-way in time and direction) Einstein-Rosen bridge is absent at all inside the Schwarzschild black hole (see for details, e.  g., [4, 6]).
@@ 5 Conclusion and Discussion
It is demonstrated the symmetry and asymmetry of the voyage on one-way Einstein-Rosen bridge inside black hole toward the endless multiplanetary future.
The apparent symmetry of both the Carter–Penrose and embedding diagrams is mainly related with a pure geometrical vision of this phenomenon. Quite the contrary, the physical (space-time) vision elucidates the absolute asymmetry of the Einstein-Rosen bridge due to existence of the light-cone limitation for possible motions.
Note, that the traversable (though only one-way in time and direction) Einstein-Rosen bridge exist only in the case of both rotating Kerr $a\neq 0$ and electrically charged Reissner–Nordström $q\neq 0$ black holes. It is absent at all inside the Schwarzschild black hole (see for details, e.  g., [4, 6]).
The infinite winding up of particle trajectories on the black hole event horizon is a manifestation of symmetry behavior of all trajectories, plunging into rotating black hole. At the same direction of the black hole rotation is a corresponding manifestation of asymmetry of the gravitational field of the Kerr metric.
We are grateful to E. O. Babichev, V. A. Berezin, Yu. N. Eroshenko, N. O. Nazarova and A. L. Smirnov for stimulating discussions.

### v4

@@ 4 Symmetry and asymmetry of test particle trajectories near rotating black hole
Figures 57 demonstrate both symmetry and asymmetry features of massive and massless particle trajectories plunging into rotating Kerr black hole with spin $a=0.9982$. Magenta arrows show the direction of the black hole rotation in accordance with the gimlet rule. Multicolored curves at Figs. 7 and 7 are the geodesic trajectories for massless test particles like photons ($\mu=0$) numerically calculated with using equations of motion (3)–(6). By approaching the black hole, the trajectories of all particles, both massive and massless ones, are infinitely winding up on the black hole event horizon in the direction of black hole rotation and at the fixed latitudes. This winding up is a symmetry manifestation of all trajectories behavior, which are plunging into rotating black hole. At the same direction of black hole rotation is a corresponding asymmetry manifestation of the gravitational field of the Kerr metric.
At last, for completeness of black hole symmetric and asymmetric properties, at Figs. LABEL:fig8 is shown the trajectory of the test planet ($\mu\neq 0$) with parameters $\gamma=1$, $\lambda=1$ and $Q=0.5$. This test planet is plunging into the spherically symmetric and nonrotating Schwarzschild black hole (with both the spin $a=0$ and electric charge $q=0$), starting from the radial distance $r=6$.
It must be especially checked that the traversable (though only one-way in time and direction) Einstein–Rosen bridge is absent at all inside the Schwarzschild black hole (see for details, e.  g., mtw ; Penrose ).
@@ 5 Conclusion and Discussion
It is demonstrated the symmetry and asymmetry of the voyage on one-way Einstein–Rosen bridge inside black hole toward the endless multiplanetary future.
The apparent symmetry of both the Carter–Penrose and embedding diagrams is mainly related with a pure geometrical vision of this phenomenon. Quite the contrary, the physical (space-time) vision elucidates the absolute asymmetry of the Einstein–Rosen bridge due to existence of the light-cone limitation for possible motions.
Note, that the traversable (though only one-way in time and direction) Einstein–Rosen bridge exist only in the case of both rotating Kerr $a\neq 0$ and electrically charged Reissner–Nordström $q\neq 0$ black holes. It is absent at all inside the Schwarzschild black hole (see for details, e.  g., mtw ; Penrose ).
The infinite winding up of particle trajectories on the black hole event horizon is a manifestation of symmetry behavior of all trajectories, plunging into rotating black hole. At the same time, the fixed direction in space of the black hole rotation axis is a strict manifestation of the Kerr metric both asymmetry and asymmetry.
We are grateful to E. O. Babichev, V. A. Berezin, Yu. N. Eroshenko, N. O. Nazarova and A. L. Smirnov for stimulating discussions. Authors also are very indebted to anonymous reviewer for suitable references and historical remarks which improve the presentation of paper.